# LEASE: Offline Preference-based Reinforcement Learning with High Sample Efficiency

### v2

@@ VI-B Results for Reward Model
VI-B1 The effect of the selection mechanism for pseudo-label accuracy
We employ a pre-trained reward model to annotate unlabeled preference data generated during the process. This paper proposes a novel screening mechanism to select data samples with both high confidence and low uncertainty, thereby enhancing the accuracy of pseudo-labels. The pseudo-label is generated by a preference model, which is constructed based on the reward model as formulated in Eq. (2). Table II compares the pseudo-label accuracy under four different conditions: 1) using the proposed screening mechanism (combining both confidence and uncertainty thresholds), 2) employing only the confidence threshold (No-U), 3) applying only the uncertainty threshold (No-C), and 4) without any screening mechanism (No-C-U). The accuracy of the pseudo-label generated by the reward model is evaluated on all preference datasets.
Table II shows that the proposed screening mechanism significantly improves pseudo-label accuracy, i.e., reduces error $\eta$ defined in Assumption 2. The pseudo-label accuracy is improved by $4.5\%$, $2.8\%$, and $6.9\%$ for pen-expert, door-expert, and hammer-expert, respectively. Moreover, we also analyze the effect of the number of preference data on the preference model. Fig. 4 illustrates the relationship between the prediction accuracy of the preference model without data augmentation and the number of preference data $N_{l}$. Here, we use the last $200$ of $2000$ preference data that are not seen in the training stage as the evaluation dataset. It indicates that as the number of preference data $N_{l}$ increases, the prediction accuracy of the preference model shows an upward trend. This also validates Theorem 1 that increasing the number of samples $N_{l}$ can reduce the generalization error.
VI-B2 The effect of data augmentation on the performance of the reward model
@@
To answer question $\mathbf{Q_{1}}$, we compare the accuracy of the reward model before and after data augmentation under a smaller preference dataset. Fig. 3 shows the comparison results between prediction and ground truth of reward, where the predicted and true rewards are both normalized to $[0,1]$. We randomly sample $500$ data points from unlabeled datasets that are not seen in the training stage for evaluation. This figure indicates that the linear relationship between reward predicted by LEASE and ground truth is better than that of the other two methods, where the method not using a selecting mechanism is denoted as FRESH. The prediction of the reward model learned by method FRESH is very narrow, and the accuracy is greatly reduced compared with FEWER. This is because the generated data of FRESH are not selected by $f(\sigma_{0},\sigma_{1})$, which causes substantial errors for the labels of generated data and leads to the collapse of training. This validates that the performance of the reward model can be improved through data augmentation and a selection mechanism, where the designed selection mechanism has a greater impact on the reward model’s performance.
@@ VI-C1 The comparison results on benchmark tasks
To answer question $\mathbf{Q_{2}}$, here, we evaluate LEASE performance based on CQL [4] and IQL [25] two offline algorithms on six types of tasks and 12 datasets in total. Table III presents a comparative analysis of he results of the average normalized score with standard deviation under three distinct scenarios: 1) performance with extensive preference data (URLHF), 2) performance with limited data without employing data augmentation techniques (FEWER), and 3) performance with limited data while incorporating data augmentation (our proposed method LEASE). In this table, m and m-e represent medium and medium-expert, respectively. The results of the offline PbRL algorithm under sufficient feedback (URLHF) are obtained from the paper [21]. As demonstrated in Table III, LEASE exhibits significantly superior performance compared to FEWER. This empirical evidence substantiates that data augmentation can effectively enhance the agent’s capability.
Furthermore, compared to the performance achieved with extensive preference data URLHF, LEASE attains comparable results in most cases. However, the proposed method exhibits slightly inferior performance on specific tasks, such as hopper-medium-expert and halfcheetah-medium-expert datasets. This limitation primarily stems from the excessively narrow distribution of the expert dataset, which consequently induces a constrained distribution in the generated preference data. As a result, the reward model’s performance cannot be effectively enhanced through data augmentation, thereby restricting the potential improvement in agent performance. However, overall, the results in Table III can validate the content of Theorem 2, which shows that LEASE can reduce the performance gap caused by the reward model and improve the agent’s performance with fewer preference data.
@@ VI-C3 The effect of the number of preference data $N_{l}$ for agent performance
Fig. 4 indicates that the number of preference data $N_{l}$ influences the accuracy of the reward model directly. URLHF [21] trains reward model with $2000$ preference dataset. This paper aims to achieve competitive performance with limited preference data, that is, achieve comparable performance with URLHF under a smaller preference dataset. We further investigate the impact of the number of preference data on the agent’s performance. Table V compares the agent’s performance when the number of preference datasets is $20$ and $100$, using the CQL algorithm for MuJoCo tasks and the IQL algorithm for Adroit tasks. This table indicates LEASE can still improve agent performance under a very small preference dataset. Notably, the performance gap between $100$ preference data and $20$ preference data is not much. This indicates that LEASE has the potential to perform well under a smaller preference dataset.
@@ VII-B Discussion for Broader Impacts
Actual application. In actual scenarios, especially for human-in-loop control, such as exoskeleton robot assistance or rehabilitation, designing a highly sample-efficient RL algorithm is greatly significant. Firstly, many rewards are difficult to describe in mathematical terms in some situations, such as the comfort of interaction. Human feedback or preference is a better way to reflect the above indices. In addition, in human-in-the-loop control, obtaining preference data through interaction can easily cause fatigue, and inappropriate interaction may cause damage to humans. It is a necessity to learn a reward model from a limited preference dataset. Then, for robot assistance or rehabilitation, the goal of the robot is to reduce human metabolic cost and guarantee the comfort of interaction. However, the reward model for this goal is difficult to design. Therefore, PbRL has a great potential to improve control performance in some scenarios where the reward function is difficult to describe and humans are in the control loop.
Theoretical study. The theory of LEASE shows that the performance gap between the behavior policy $\mu$ and $\widehat{\pi}$ learned by LEASE includes two part: the gap $\xi$ caused by offline algorithm itself and the gap $\xi_{1}$ caused by reward model gap (the details see Theorem 2), where the term $\xi_{1}$ only depends on preference dataset. Moreover, the theory of LEASE is based on state-action pairs, which is consistent with most offline algorithms theory. Therefore, the theory developed in this paper can be easily combined with other offline algorithm theories and easily used to build the theory of policy improvement guarantee, which can provide the theoretical basis for offline PbRL and facilitate the development of offline PbRL theory.
@@ VIII Conclusion
This paper proposes a novel offline PbRL algorithm (LEASE) with high sample efficiency. LEASE can achieve comparable performance under fewer preference datasets. By selecting high-confidence and low-variance data, the stability and accuracy of the reward model are guaranteed. Moreover, this paper provides the theoretical analysis for LEASE, including the generalization of the reward model and policy improvement guarantee. The theoretical and experimental results demonstrate that the data selection mechanism $f(\sigma_{0},\sigma_{1})$ can effectively improve the performance of the reward model, and the performance learned by LEASE can be guaranteed under fewer preference datasets. Future work can focus on reducing this gap further under limited preference data or on achieving conservative estimation for state-action pairs where the learned reward model predicts inaccurately.
@@ A-B Proof of Theorem 1
The Proof for Theorem 1: This proof is based on Proposition 1 and 2. Combing Eqs. (A.3) and (A.4), we can derive that for labeled dataset $\widehat{\mathcal{D}}=\{(\sigma_{0},\sigma_{1},y)^{(i)}\}_{i=1}^{N}$, the generalization error bound between expected error $\mathcal{L}(\psi)$ and empirical error $\widehat{\mathcal{L}}(\psi)$ holds:
@@ A-E1 The other comparison results for reward model performance
Fig. A.1 shows the comparison between the prediction value by the learned rewards and their ground truths for LEASE, FEWER, and FRESH under Adroit tasks, where the offline algorithm is IQL. The predicted and true rewards are both normalized to $[0,1]$. We randomly sample $500$ data from unlabeled datasets that are not seen in the training stage for evaluation, like Fig. 4. Since the reward has a value in certain positions in Adroit tasks, the figure is in the form of multiple scattered lines in Fig. A.1. As evidenced by the experimental results shown in this figure, the reward model trained under the LEASE framework achieves the smallest variance in predicted reward values, which shows that the data augmentation and selection mechanism can improve the reward model’s performance.
@@ A-E3 The results for LEASE under model-based RL offline algorithm
To enhance the generalizability of the proposed framework, we extended our experimental validation beyond model-free offline RL approaches (such as CQL and IQL) to include model-based offline RL methodologies. Table A.2 shows the results of COMBO [49] under the designed framework, where * denotes the results of offline RL algorithms under our designed framework. For COMBO hyperparameters, the probability confidence $\kappa_{p}$, and uncertainty variance $\kappa_{\tau}$ are set as $0.85$ and $0.05$ for all tasks, respectively. Note that in our framework, model-based methods do not necessarily perform better than model-free methods. Model-based RL methods focus on how to learn a conservative policy by regularizing $Q$ values or penalizing rewards to alleviate the effects of inaccurate model data. These methods critically depend on the accuracy of the reward model. Consequently, model-based approaches are more sensitive to reward modeling errors than model-free methods, which can hinder the improvement of agent performance. Therefore, model-based offline RL requires a higher accuracy of the reward model than model-free RL.

### v3

@@ VI-B Results for Reward Model
VI-B1 The effect of data augmentation on the performance of the reward model
@@
To answer question $\mathbf{Q_{1}}$, we compare the accuracy of the reward model before and after data augmentation under a smaller preference dataset. Fig. 3 shows the comparison results between prediction and ground truth of reward, where the predicted and true rewards are both normalized to $[0,1]$. We randomly sample $500$ data points from unlabeled datasets that are not seen in the training stage for evaluation. This figure indicates that the linear relationship between reward predicted by LEASE and ground truth is better than that of the other two methods, where the method not using a selecting mechanism is denoted as FRESH. The reward model learned by the method FRESH exhibits a notably narrow prediction range, and its accuracy significantly drops compared with that of FEWER. This is because the data generated by FRESH are not filtered by $f(\sigma_{0},\sigma_{1})$, leading to substantial errors in data labeling and consequently causing training collapse. These results validate that the performance of a reward model can be enhanced through data augmentation and selection mechanisms, with the design of the selection mechanism having a more pronounced impact on the model’s overall performance.
TABLE III: Comparison results of pseudo-labeling accuracy between using the selection mechanism (Ours), only using confidence (No-U), only using uncertainty (No-C), and not using the selection mechanism (No-C-U).
Task Name Ours No-U No-C No-C-U
pen-expert $\mathbf{87.3\pm 0.5}$ $85.9\pm 0.8$ $86.7\pm 0.6$ $86.8\pm 0.5$
door-expert $\mathbf{89.3\pm 0.6}$ $87.8\pm 0.2$ $86.0\pm 0.0$ $85.9\pm 0.8$
Average $\mathbf{87.4\pm 0.4}$ $86.0\pm 0.5$} $85.1\pm 0.6$} $85.4\pm 0.6$}
VI-B2 The effect of the selection mechanism for pseudo-label accuracy
This paper introduces a novel screening mechanism designed to select unlabeled preference data samples characterized by both high confidence and low uncertainty, thereby improving pseudo-label accuracy. The pseudo-labels are generated by a preference model derived from a pre-trained reward model, as defined in Eq. (2). The performance of this approach is evaluated in terms of pseudo-label accuracy across all preference datasets. As shown in Table III, the proposed screening mechanism—which simultaneously applies confidence and uncertainty thresholds—is compared against three baseline conditions: employing only the confidence threshold (No-U), employing only the uncertainty threshold (No-C), and using no screening mechanism (No-C-U).
Table III demonstrates that the proposed screening mechanism substantially enhances the pseudo-label accuracy, i.e., it reduces the error $\eta$ defined in Assumption 2. For the Pen, Door, and Hammer experts, the pseudo-label accuracy is improved by 4.5%, 2.8%, and 6.9%, respectively. Furthermore, we analyze the influence of the amount of preference data on the preference model. Fig. 4 illustrates the relationship between the prediction accuracy of the preference model—without data augmentation—and the quantity of preference data $N_{l}$. Here, we use the final 200 preference samples (out of a total of 2000) that were not observed during training as the evaluation dataset. The results indicate that as the amount of preference data $N_{l}$ increases, the prediction accuracy of the preference model exhibits an upward trend. This observation also corroborates Theorem 1, which states that increasing the number of samples $N_{l}$ reduces the generalization error.
@@ VI-C1 The comparison results on benchmark tasks
To answer question $\mathbf{Q_{2}}$, here, we evaluate LEASE performance based on CQL [4] and IQL [25] two offline algorithms on six types of tasks and 12 datasets in total. Table II presents a comparative analysis of he results of the average normalized score with standard deviation under three distinct scenarios: 1) performance with extensive preference data (URLHF), 2) performance with limited data without employing data augmentation techniques (FEWER), and 3) performance with limited data while incorporating data augmentation (our proposed method LEASE). In this table, m and m-e represent medium and medium-expert, respectively. The results of the offline PbRL algorithm under sufficient feedback (URLHF) are obtained from the paper [21]. As demonstrated in Table II, LEASE exhibits significantly superior performance compared to FEWER. This empirical evidence substantiates that data augmentation can effectively enhance the agent’s capability.
Furthermore, compared to the performance achieved with extensive preference data URLHF, LEASE attains comparable results in most cases. However, the proposed method exhibits slightly inferior performance on certain specific tasks, such as the hopper-medium-expert and halfcheetah-medium-expert datasets. This limitation primarily stems from the overly narrow distribution of the expert datasets, which consequently constrains the distribution of the generated preference data. As a result, data augmentation fails to effectively enhance the performance of the reward model, thereby limiting the potential improvement of the agent’s performance. However, overall, the results in Table II can validate the content of Theorem 2, which shows that LEASE can reduce the performance gap caused by the reward model and improve the agent’s performance with fewer preference data.
@@ VI-C3 The effect of the number of preference data $N_{l}$ for agent performance
Fig. 4 indicates that the number of preference data $N_{l}$ influences the accuracy of the reward model directly. URLHF [21] trains reward model with $2000$ preference dataset. This paper aims to achieve competitive performance with limited preference data, that is, achieve comparable performance with URLHF under a smaller preference dataset. We further investigate the impact of the number of preference data on the agent’s performance. Table V compares the agent’s performance when the number of preference datasets is $20$ and $100$, using the CQL algorithm for MuJoCo tasks and the IQL algorithm for Adroit tasks. This table indicates LEASE can still improve agent performance under a very small preference dataset. It is noteworthy that the performance gap between using 100 preference datasets and 20 preference datasets is not substantial. This is primarily because the rewards in the D4RL dataset are relatively simple; a small amount of preference data is sufficient to learn a reasonably accurate reward model. Coupled with data augmentation, the performance of the reward function can be further improved, thereby enhancing the agent’s overall capability. This indicates that LEASE has the potential to perform well even with a limited number of preference datasets.
@@ VII-B Discussion for Broader Impacts
Actual application. The high-sample-efficiency preference-based offline RL algorithm proposed in this work aims to learn robot control policies from offline data with minimal human feedback, under the condition of an unknown reward function. This approach is particularly suitable for scenarios where the reward function is difficult to specify mathematically and where human-robot interaction entails high cost. For instance, in the field of exoskeleton robot assistance, objectives such as maximizing assistive comfort and minimizing metabolic cost are challenging to define precisely, and prolonged human-in-the-loop interaction can lead to user fatigue. Therefore, it is imperative to learn a reward model from a limited preference dataset and utilize it to efficiently guide policy optimization.
The deployment of the algorithm for practical control involves three main stages: 1) collecting an offline dataset (containing only $(s,a,s^{\prime})$ tuples) through methods such as teleoperation, to train a transition model; 2) acquiring a limited set of preference data by having users select their preference between pairs of trajectories, which is then used to learn a reward function; and 3) training the robot control policy offline using the collected dataset and the learned reward model, followed by its deployment to the actual control system. The ultimate goal is to optimize a control policy offline with minimal human-robot interaction efficiently.
Theoretical study. The theory of LEASE shows that the performance gap between the behavior policy $\mu$ and $\widehat{\pi}$ learned by LEASE includes two part: the gap $\xi$ caused by offline algorithm itself and the gap $\xi_{1}$ caused by reward model gap (the details see Theorem 2), where the term $\xi_{1}$ only depends on preference dataset. Moreover, the theory of LEASE is based on state-action pairs, which is consistent with most offline algorithms theory. The performance gap in offline PbRL is determined by two independent components, with the first term corresponding to the inherent gap of offline RL itself. Given this gap, the preference-based gap can be directly derived. Therefore, the theory developed in this paper can be easily combined with other offline algorithm theories and used to build a theory of policy improvement guarantee, providing a theoretical basis for offline PbRL and facilitating the development of offline PbRL theory.
@@ VIII Conclusion
This paper proposes a novel offline PbRL algorithm (LEASE) with high sample efficiency. LEASE can achieve comparable performance under fewer preference datasets. By selecting high-confidence and low-variance data, the stability and accuracy of the reward model are guaranteed. Moreover, this paper provides the theoretical analysis for LEASE, including the generalization of the reward model and policy improvement guarantee. The theoretical and experimental results demonstrate that the data selection mechanism $f(\sigma_{0},\sigma_{1})$ can effectively improve the performance of the reward model, and the performance learned by LEASE can be guaranteed under fewer preference datasets. Future works can focus on the following two aspects: 1) Since the trained transition model may have errors, which will affect the performance of the reward model to some extent, the quality of generated data can be considered in the future to further improve the accuracy of the reward function; 2) There is still a gap between the trained reward model and the real model. Other efficient methods of data augmentation can be used in the future to further narrow this gap.
@@ A-B Proof of Theorem 1
The Proof for Theorem 1: This proof is based on Proposition 1 and 2. Combining Eqs. (A.3) and (A.4), we can derive that for labeled dataset $\widehat{\mathcal{D}}=\{(\sigma_{0},\sigma_{1},y)^{(i)}\}_{i=1}^{N}$, the generalization error bound between expected error $\mathcal{L}(\psi)$ and empirical error $\widehat{\mathcal{L}}(\psi)$ holds:
@@ A-E1 The other comparison results for reward model performance
Fig. A.1 shows the comparison between the prediction value by the learned rewards and their ground truths for LEASE, FEWER, and FRESH under Adroit tasks, where the offline algorithm is IQL. For evaluation, we randomly sample $500$ data points from the unlabeled dataset that are not seen during the training phase, as illustrated in Fig. 4. Because rewards in Adroit tasks are limited to a few specific values (i.e., discrete states), while rewards in Mujoco tasks are continuous, Fig. A.1 is presented as multiple scatter lines, in contrast to Fig. 4. As evidenced by the experimental results shown in this figure, the reward model trained under the LEASE framework achieves the smallest variance in predicted reward values, which shows that the data augmentation and selection mechanism can improve the reward model’s performance.
@@ A-E3 The results for LEASE under model-based RL offline algorithm
To enhance the generalizability of the proposed framework, we extended our experimental validation beyond model-free offline RL approaches (such as CQL and IQL) to include model-based offline RL methodologies. Table A.2 shows the results of COMBO [49] under the designed framework, where * denotes the results of offline RL algorithms under our designed framework. For COMBO hyperparameters, the probability confidence $\kappa_{p}$, and uncertainty variance $\kappa_{\tau}$ are set as $0.85$ and $0.05$ for all tasks, respectively. It is important to note that, within our framework, model-based methods do not necessarily outperform model-free approaches. Model-based offline reinforcement learning typically emphasizes learning conservative policies through techniques such as regularizing Q-values or penalizing rewards, which mitigates the impact of inaccuracies in the model-generated data. However, these methods rely heavily on the accuracy of the reward model. Consequently, model-based approaches are more sensitive to errors in the reward model than their model-free counterparts, which may constrain potential improvements in agent performance. Therefore, model-based offline reinforcement learning generally requires a higher degree of reward model accuracy compared to model-free methods.